OpenStudy (anonymous):

The Lewiston Landscaping Company charged Mr. Rabinowitz $100 for mowing his rectangular lawn, which is 20 feet long and 12 feet wide. If Mr. Nabors's lawn is also rectangular and measures 15 feet long and 10 feet wide, and the price charged for mowing a lawn is proportional to its area, which of the following expressions can be used to calculate how much (x) the company will charge Mr. Nabors?

OpenStudy (anonymous):

first yard has area \[12\times 20=240\] square feet second one has area \[15\times 10=150\] square feet. solve the proportion \[\frac{100}{240}=\frac{x}{150}\] \[x=\frac{150\times 100}{240}\]